OpenStudy (oblivion20):

Gina called five electronics stores to find their price for a certain cell phone charger. She recorded the results in a table. What was the mean price of the cell phone chargers? Store Price ($) Best Deals 31 Electronics Hut 21 Music Unlimited 26 All Things Digital 23 Wireless Haven 19 medal if helped

OpenStudy (automaticloveletter):

31+21+26+23+19=120/5=24 How many companies are there? 5, that is why we divide by five to find the average or "mean price" Why are we adding? Add all of the prices of each phone company to find the total More help understanding mean at: http://www.purplemath.com/modules/meanmode.htm

OpenStudy (oblivion20):

thank you @automaticloveletter you're awesome

OpenStudy (automaticloveletter):

Thanks :))